OpenStudy (anonymous):

Simplify 7x2 + 4.5y – 3y + x2 – 5x The answer is 8x^2 -5x + 1.5y Please explain how someone got this answer, thanks.

OpenStudy (anonymous):

combine like terms first

OpenStudy (anonymous):

4.5 y - 3.5 y = 1.5 y like this?

OpenStudy (anonymous):

7x^2 + x^2 = 8^2

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

all terms with x^2 are like terms, all with x are like terms, and all with y are like terms

OpenStudy (anonymous):

you already got the x^2's so lets do the x's: theres only one x which is -5 so we keep it like that

OpenStudy (anonymous):

now y's :theres 4.5 and -3 so 4.5 - 3 = 1.5. the two y terms combined = 1.5y

OpenStudy (anonymous):

I get it now, Thanks!

OpenStudy (anonymous):

+1

OpenStudy (anonymous):

anytime